Mars Global Surveyor Update
18 Sep 2002
(Source: Jet Propulsion Laboratory)
Mars Global Surveyor Status Report
Wednesday, September 18, 2002 (DOY 254/19:00:00 to DOY 261/19:00:00 UTC)
Launch / Days since Launch = Nov. 7, 1996 / 2142 days
Start of Mapping / Days since Start of Mapping = April 1, 1999 / 1267 days
Total Mapping Orbits = 15,779
Total Orbits = 17,462
Background Sequences - The spacecraft is operating nominally in performing the beta-supplement daily recording and transmission of science data. The mm209 sequence executed successfully from 02-255 (9/12/02) through 02-257 (9/14/02). The mm210 sequence has performed well since it started on 02-258 (9/15/02). It terminates on 02-261 (9/18/02). The mm211 sequence, successfully uplinked on 02-260 (9/17/02), begins executing near the end of day on 02-261 (9/18/02).
Other - The three Roll Only Targeted Observations (ROTOs) reported as successful last week were in fact not successful. The images taken during those ROTOs and the first two ROTOs in mz200 were distorted because the spacecraft was slewing slowly while the images were being taken. The relative slew rate parameter had been set to orbital rate while the equator-crossing software patch was being loaded on 02-248 (9/05/02). The parameter was not reset to zero following the patch update. The problem was discovered on 02-255 (9/12/02) and corrected on 02-256 (9/13/02). The following four ROTOs in mz200 and six ROTOs in mz201 were performed successfully. MGS has successfully completed 306 ROTOs to date.
Spacecraft subsystems report good health and status.
There have been 23 uplinks to the spacecraft during the past week, including new star catalog and ephemeris files; instrument command loads; the mi0654 slew-rate parameter update; the mz200, mz201 & mz202 ROTO mini-sequences; and the mm210 & mm211 background sequences. 7,040 command files have been radiated to the spacecraft since launch.
ROTOs are planned for next week in the mz201, mz202, & mz203 mini-sequences. Radio Science Occultation Egress Scans will be performed on 02-273 (9/30/02) and 02-274 (10/01/02). A DDOR experiment is scheduled for 02-276 (10/03/03).